Multiple Choice Question 74

Waterway Industries issues 4000 shares of its $5 par value common stock having a fair value of $25 per share and 6000 shares of its $10 par value preferred stock having a fair value of $25 per share for a lump sum of $205100. What amount of the proceeds should be allocated to the preferred stock?

$139367
$123060
$82039
$178969

Multiple Choice Question 90

Sheffield Corp. owned 15100 shares of Pharoah Company. These shares were purchased in 2014 for $105700. On November 15, 2018, Sheffield declared a property dividend of one share of Pharoah for every ten shares of Sheffield held by a stockholder. On that date, when the market price of Pharoah was $29 per share, there were 105700 shares of Sheffield outstanding. What gain and net reduction in retained earnings would result from this property dividend?

Gain

Net Reduction in
Retained Earnings

$0          

$73990

$232540

$73990

$232540

$ 28000

$0          

$306530
Answer #1

74.

Fair values of common stock and preferred stock

= (4000*25) + (6000*25) = 250,000

Preferred stock's percentage = (6000*25)/250,000 = 60%

Proceeds allocated to preferred shars

= 205,100 * 60%

= 123,060

Option B